Nation’s Capital Fitness, Inc. operates a chain of fitness centers in the Washington, D.C., area. The firm’s controller is accumulating data to be used in preparing its annual profit plan for the coming year. The cost behavior pattern of the firm’s equipment maintenance costs must be determined. The accounting staff has suggested the use of the high-low method to develop an equation, in the form of Y = a + bX, for maintenance costs. Data regarding the maintenance hours and costs for last year are as follows:

Month

Hours of Maintenance
Service

Maintenance
Costs

January

530

$

4,975

February

480

4,220

March

280

2,750

April

450

4,290

May

340

3,050

June

430

4,180

July

340

3,070

August

450

3,540

September

490

4,030

October

350

3,250

November

350

3,210

December

330

3,070

Total

4,820

$

43,635

Average

402

*

$

3,636

*

*Rounded
Answer #1

Using the high-low method of cost estimation, estimate the behavior of the maintenance costs incurred by Nation’s Capital Fitness, Inc.

Hours Cost
High January      530.00 $ 4,975.00
Low March      280.00 $ 2,750.00
$ 250.00 $ 2,225.00

Variable Cost Per Hour = 2,225/250 = 8.9

Fixed Cost = 4,975 - (8.9 x 530) = $258

Equation in y = a + bx
4,975 = $258 + (8.9 x 530)